OpenStudy (anonymous):

What is the corrent way to divide 2x3 + 6x2 - 2x - 5 by x + 3 using synthetic division? The 3 and two are upper ones

OpenStudy (misty1212):

i hope you have choices

OpenStudy (nincompoop):

have you got any clue what synthetic division is?

OpenStudy (misty1212):

no

OpenStudy (misty1212):

jk

OpenStudy (nincompoop):

because you don't need choices if you know what it is and how to do it

OpenStudy (misty1212):

because your problem says "what is the correct way" i am guessing you have some choices otherwise it would just probably say "divide"

OpenStudy (misty1212):

yikes looks like a gahan wilson cartoon

OpenStudy (anonymous):

A.. -3 2 6 -2 -5 -6 0 6 = 20-21 B. 3 2 6 -2 -5 6 36 102 = 2 12 34 97

OpenStudy (anonymous):

theres a and b

OpenStudy (misty1212):

ok skip B there should be a -3 in the box up top,not a 3

OpenStudy (anonymous):

okay if its not a it has to be c then

OpenStudy (anonymous):

only other one with -3

OpenStudy (misty1212):

probably is the bottom line \[2,0,2,1\] for C?